OpenStudy (anonymous):

solve the equation (x-2)(3x+8)=0

OpenStudy (anonymous):

if you have two factors, what do you need to have their product equal to zero?

OpenStudy (anonymous):

?????

OpenStudy (unklerhaukus):

at least one of the factors must be zero

OpenStudy (anonymous):

that is right, how do you apply this to this product?

OpenStudy (anonymous):

ok...i have to say it but i understand one factor must be zero. After that I 'm lost

OpenStudy (anonymous):

can you write your factors separately?

OpenStudy (anonymous):

ill give you a clue. one of your factors is x-2, what is the other?

OpenStudy (anonymous):

3x+8

OpenStudy (anonymous):

thats right. in order to have their product equal to zero, x-2=0 or 3x+8=0

OpenStudy (anonymous):

can you solve that?

OpenStudy (anonymous):

x = 2 and 8/3???

OpenStudy (anonymous):

sorry -8/3

OpenStudy (anonymous):

careful! when you move the 8 to the right ef sign equal , something happens with its sign, right?

OpenStudy (anonymous):

thats much better (some telepathy between you and I)

OpenStudy (anonymous):

@MixMasterTX have you got the idea?

OpenStudy (anonymous):

yes @CarlosGP x = 2, -8/3 Thanks!

OpenStudy (anonymous):

u r welcome

OpenStudy (unklerhaukus):

\[\large\color{red}\checkmark\]

OpenStudy (anonymous):

@UnkleRhaukus thank you for your helping hand too